OpenStudy (anonymous):

$7.45 an hour and worked 47 hours this week what is the gross pay

OpenStudy (anonymous):

350.15 I hope this is your answer!

OpenStudy (anonymous):

7.45 x 40 = 298 if there is time and a half for overtime... 7.45 x .5(7.45) = 7.45 + 3.725 = 11.18 per overtime hr. 7 x 11.18 = 78.26 regular pay + overtime pay = gross pay 298 + 78.26 = 376.26